4-Kg. 5 kg (C) 4 kg string. The disc acts like a 22 olid of mass 16kg and radius 50 cm has a mass of4 kg hanging from one side by a massless torotate disc center. Once released from rest, how much time does it take for the wheel and rotates A.uhrnuts thech through 20 revolutions? 22 seconds (A) seconds (c) 2vir seconds. seconds 4-kg
Answer #1

By using Free Body Diagram we can write equilibrium equations

4g-T = 4aFor Block of 4Kg

Where T is tension in the string and a is acceleration downwards.

Similarly, For disc,

4g-T = 4a

As there is no slipping a = r\alpha

So, Equation for disc will be

Tr = \frac{1}{2}16r^2*\frac{a}{r}Tr = I\alpha

T = 8a

Use this valye of T in equation og block

4g = 12a

a = \frac{g}{3}

Diatnce travelled by block = 20*2\pi*0.5 = 20\pi

By equation of motion

S = ut + \frac{1}{2}at^2

20\pi = \frac{1}{2}*\frac{10}{3}t^2

t = 2 \sqrt{3\pi} sec

This is not given in the option.